solve each absolute value equation.
3|w+1|-2=7

Answer by stanbon(57342)

3|w+1|-2=7
---
3|w+1|= 9
------
|w+1| = 3
---
w+1 = 3 or w+1 = -3
w = 2 or w = -4
================
Cheers,
Stan H.
=============